Please answer question 2.17

jonnsor Re8O 216 Directjet is planning to purchase two Airbus A350-1000 aircraft at the price of 320 million each in eight years. Directjet currently has e320 million in an investment account. At approximately what compound interest rate can it accumulate 640 million in eight years to acquire the two A350-1000 aircraft? 2.17 In Problem 2.16, assume all relevant data, at approximately what simple interest rate can DirectJet accumulate 640 million in eight years to acquire the two A350-1000 aircraft?
Answer #1

2.17) Currently DirectJet has 320 million. In 8 years the company wants to accumulate 640 million. It means it has to earn 320 million as interest amount. If we divide this 320 million equally in 8 years, per year interest amount = 320 / 8 = 40 million

The Simple interest rate = (40 / 320) * 100 = 12.5%

Thus, at a simple interest rate of 12.5% DirectJet can accumulate 640 million in 8 years to acquire the two A350-1000 aircraft.